OpenStudy (anonymous):

find the fourth term of the geometric sequence whose first term is 5 and whose common ratio is -2

OpenStudy (anonymous):

5+[(4-1)*(-2)] 4+[3*(-2)]= 3+(-6) = -3

OpenStudy (anonymous):

thank u..

OpenStudy (anonymous):

uwc